OpenStudy (anonymous):

a vase is formed by rotating the portion of the curve y^2= x-6 between y=6 and y=-6

OpenStudy (anonymous):

about the y axis. Find the volume of the vase

OpenStudy (anonymous):

Oops. Typo, hold on..

OpenStudy (anonymous):

This is a disk method problem, which I assume you are familiar with. The radius of each disk is x = y^2 + 6 So the area of each disk is pi * (y^2+6)^2 = pi (y^4 + 72y^2 + 36) and the volume of each infinitesimal slice is pi (y^4 + 72y^2 + 36) dy Thus we want to evaluate the following integral: \[V = \pi \int_{-6}^{6} y^4 + 72 y^2 + 36 \space dy = 2\pi \int_0^6 y^4 + 72 y^2 + 36 \space dy\]

OpenStudy (anonymous):

Oh damn it i did it again. that's not 72, it's 12 :)

OpenStudy (anonymous):

thank you very much jemurray3!